OpenStudy (freemap):

The width of a rectangle is 8 cm less than its length. The perimeter of the rectangle is 52 cm. What is the width of the rectangle? A. 7 cm B. 9 cm C.15 cm D. 17cm

OpenStudy (freemap):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`The width of a rectangle is 8 cm less than its length` if the length is L, then what is the width in terms of L?

OpenStudy (freemap):

W=L-8

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now we use the perimeter of a rectangle formula P = 2L + 2W

jimthompson5910 (jim_thompson5910):

replace P with 52 replace W with L-8 doing this gives 52 = 2L + 2(L-8) now solve for L

OpenStudy (freemap):

Would you divide

jimthompson5910 (jim_thompson5910):

yes but not right away

OpenStudy (freemap):

l'm sorry but what would be L. Like the number I really don't understand

jimthompson5910 (jim_thompson5910):

L stands for the length it's unknown at this point, but when you solve for L, you figure out what the number for the length is

jimthompson5910 (jim_thompson5910):

example: if 2L = 10, then L = 5

OpenStudy (freemap):

52-2L =26

jimthompson5910 (jim_thompson5910):

what would 2(L-8) turn into?

OpenStudy (freemap):

2L -8

OpenStudy (freemap):

2L - 16 i mean

jimthompson5910 (jim_thompson5910):

so 52 = 2L + 2(L-8) turns into 52 = 2L + 2L-16

jimthompson5910 (jim_thompson5910):

then you combine the 2L terms 2L+2L = 4L making 52 = 2L + 2L-16 turn into 52 = 4L-16

OpenStudy (freemap):

Which now you divide by by 4 now?

jimthompson5910 (jim_thompson5910):

not yet

OpenStudy (freemap):

oh 16 + 52 is 68 and 68 divided by 4 is 17

jimthompson5910 (jim_thompson5910):

yep, L = 17

OpenStudy (freemap):

Thanks so much much could be check my other answer i'll post again so you ca get other medal

jimthompson5910 (jim_thompson5910):

since L = 17, the width is W = L-8 = 17-8 = 9

OpenStudy (freemap):

Thanks, i forgot about that part

jimthompson5910 (jim_thompson5910):

they put that trap in